Consider the acid dissociation behavior of carboni
Answer #1


(1) Since pKa1 (6.351) < pH (8.70) < pKa2 (10.329)

=> predominant species at pH 8.70 is HCO3-


(2) pH = pKa2 + log([CO32-]/[HCO3-])

9.90 = 10.329 + log([CO32-]/[HCO3-])

log([CO32-]/[HCO3-]) = -0.429


[CO32-]/[HCO3-] = 10^(-0.429) = 0.372